# A Survey of AI Scientists

### v4

@@ 1 Introduction
Artificial intelligence (AI) is undergoing a fundamental transformation in its role within scientific research. Historically deployed as a computational instrument for specific analytical tasks (such as pattern recognition, data mining, and predictive modeling [liu2022real, almarie2023editorial, erduran2024impact]), AI is now emerging as an autonomous originator of scientific knowledge. This shift represents more than incremental technological progress; it embodies a qualitative leap from AI for Science, where AI augments human-defined workflows, to AI as Scientist, where AI systems independently conceptualize, execute, and communicate original research. The AI scientist paradigm pursues end-to-end autonomous discovery—executing the complete scientific workflow from hypothesis formulation through experimental execution to manuscript generation [gridach2025agentic, wei2025ai, Yamada2025AIScientistV2]. This transformation is enabled by three converging technological advances: (1) large language models with emergent reasoning capabilities [OpenAI2024GPT4]; (2) multi-agent architectures enabling complex task decomposition and collaboration [schmidt2009distilling, Ghafarollahi2024SciAgents]; and (3) robotic experimentation platforms bridging digital reasoning with physical execution [Curie2025, BioPlanner2023]. Collectively, these advances mark a transition from automation—wherein AI supports predefined protocols—to autonomy—wherein AI independently designs, adapts, and validates investigative processes.
The intellectual lineage of autonomous scientific discovery extends to early symbolic AI systems [Langley1987ScientificDiscovery, schmidt2009distilling], which automated rediscovery of established physical laws through symbolic regression and rule-based reasoning. However, these pioneering efforts were constrained by brittle knowledge representations, limited generalization beyond carefully curated domains, and inability to interface with real-world experimentation. Contemporary AI Scientist systems transcend these limitations by integrating foundation models with closed-loop scientific reasoning (observe → hypothesize → experiment → analyze → publish), through agentic planning, principle-driven inference, and adaptive self-correction mechanisms. This integration enables systems to operate across heterogeneous knowledge modalities, navigate ambiguous problem spaces, and iteratively refine hypotheses based on empirical feedback. Representative implementations [Lu2024AIScientist, Curie2025, PiFlow2025] demonstrate cross-domain capabilities spanning chemical synthesis [Boiko2023Coscientist], genomic network inference [Afonja2024LLM4GRN], and symbolic equation discovery [SRScientist2025], validating the generality of the autonomous research paradigm.
@@
Despite these advances, the field faces a critical challenge: conceptual fragmentation. The rapid proliferation of AI Scientist systems—evidenced by specialized frameworks achieving near-human performance [Yamada2025AIScientistV2, Tang2025AIResearcher] and domain-specific benchmarks evaluating novelty, causality, and reproducibility [chen2025auto, liu2025researchbench, guo2025ideabench]—has outpaced the development of unifying theoretical frameworks. Current research remains fragmented across three dimensions: (1) methodological silos, where systems excel at isolated stages (e.g., hypothesis generation [Zhou2024HypothesisGenerationLLMs, vasu2025hyper], literature synthesis [Agarwal2024LitLLM], or experimental execution [Curie2025]) without addressing end-to-end integration; (2) domain-specific architectures, optimized for chemistry [Boiko2023Coscientist], biology [BioPlanner2023], or physics with limited cross-domain generalization; and (3) evaluation heterogeneity, where benchmarks assess non-overlapping capabilities without establishing unified quality metrics. While existing surveys [gridach2025agentic, wei2025ai, luo2025llm4sr] provide valuable domain-specific or capability-focused reviews, no comprehensive synthesis systematically connects methodological stages, architectural patterns, and evaluation protocols across the full spectrum of AI Scientist research. This conceptual gap impedes cumulative progress, obscures transferable insights, and hinders the identification of fundamental research challenges versus domain-specific engineering issues.
This survey synthesizes research from 2022–2025 through a unified architectural framework (Figure 1). The horizontal axis defines six methodological stages: (1) Literature Review: knowledge structuring [Agarwal2024LitLLM]; (2) Idea Generation: hypothesis formulation [vasu2025hyper]; (3) Experimental Preparation: protocol design [Lai2023DS1000, guo2024ds]; (4) Experimental Execution: adaptive orchestration [BioPlanner2023, Curie2025, Panapitiya2025AutoLabs]; (5) Scientific Writing: multimodal composition [WritingBench2025Benchmark]; and (6) Paper Generation: manuscript synthesis [Tang2025AIResearcher, Lu2024AIScientist]. The vertical axis stratifies four abstraction layers—Domains, Tasks, Architectures, Capabilities—forming a 4×6 categorization matrix. The timeline panel traces three evolutionary phases: Foundational Modules, Closed-Loop Integration, and Scalable Collaboration.
@@ 2 Background and Taxonomy
DS-1000 [Lai2023DS1000] 2023.04
Coscientist [Boiko2023Coscientist] 2023.06
BioPlanner [BioPlanner2023] 2023.10
MLAgentBench [huang2023mlagentbench] 2023.10
LitLLM [Agarwal2024LitLLM] 2024.02
AI Scientist (v1) [Lu2024AIScientist] 2024.08
SciAgents [Ghafarollahi2024SciAgents] 2024.09
IdeaBench [guo2025ideabench] 2024.11
Quantum-Agent-SDL [Cao2024QuantumAgentSDL] 2024.12
HypER [vasu2025hyper] 2025.01
AI Scientist (v2) [Yamada2025AIScientistV2] 2025.02
Curie [Curie2025] 2025.02
AI co-scientist [gottweis2025towards] 2025.02
ResearchBench [liu2025researchbench] 2025.03
DeepResearcher [Zheng2025DeepResearcher] 2025.04
AutoLabs [Panapitiya2025AutoLabs] 2025.04
AI-Researcher [Tang2025AIResearcher] 2025.05
EXP-Bench [Kon2025EXPBench] 2025.05
Agentic AutoSurvey [liu2025agentic] 2025.09
PiFlow [PiFlow2025] 2025.09
DeepScientist [Weng2025DeepScientist] 2025.09
SR-Scientist [SRScientist2025] 2025.10
Freephdlabor [Li2025FreephdLabor] 2025.10
@@ 2.1 Taxonomy of Research Categories
    Literature Review (Lit.). This foundational stage involves transforming unstructured scientific corpora into machine-interpretable knowledge. It encompasses techniques from large-scale information retrieval to the synthesis of research gaps. Systems like LitLLM [Agarwal2024LitLLM] and HypER [vasu2025hyper] focus on citation-grounded summarization and knowledge extraction, while advanced frameworks like SciAgents [Ghafarollahi2024SciAgents] and DeepResearcher [Zheng2025DeepResearcher] utilize web-scale interactions and graph reasoning to map existing knowledge and provide a robust foundation for downstream tasks.
@@
    Idea Generation (Idea). Building upon the structured knowledge from the prior stage, this phase automates hypothesis discovery and problem formulation. It leverages the creative and reasoning capabilities of LLMs to propose novel yet plausible research directions. This capability is explicitly evaluated by benchmarks like IdeaBench [guo2025ideabench], and is a core component of both domain-specific systems like Coscientist [Boiko2023Coscientist] and advanced end-to-end frameworks like DeepScientist [Weng2025DeepScientist].
@@
    Experimental Preparation (Exp.). This crucial intermediate stage translates an abstract hypothesis into an executable plan. It includes tasks such as defining variables, selecting datasets, generating analysis code, and designing experimental protocols. This is a primary focus of data-science-oriented benchmarks like DS-1000 [Lai2023DS1000] and MLAgentBench [huang2023mlagentbench], and is a key step in all integrated systems from The AI Scientist v1 [Lu2024AIScientist] to freephdlabor [Li2025FreephdLabor].
@@
    Experimental Execution (Exec.). This stage involves the actual running of real or simulated experiments. It emphasizes the agent’s ability to interact with tools, control robotics, and adapt its plan based on real-time feedback. Milestones in this area include systems that orchestrate physical laboratory hardware, such as Coscientist [Boiko2023Coscientist] and Quantum-Agent-SDL [Cao2024QuantumAgentSDL]. Frameworks like Curie [Curie2025] and DeepResearcher [Zheng2025DeepResearcher] demonstrate this capability in simulated and real-world web environments, respectively.
@@
    Scientific Writing (Writ.). This stage focuses on the communication of scientific findings by transforming structured results into coherent, citation-grounded narratives. Capabilities range from section-aware summarization to data-to-text synthesis. This is a key feature in end-to-end systems like ResearchBench [liu2025researchbench] and is central to human-in-the-loop frameworks like freephdlabor [Li2025FreephdLabor], where the AI drafts content for human review and refinement.
@@
    Paper Generation (Paper). Representing the culmination of the scientific workflow, this final stage synthesizes a full, publication-ready manuscript. This requires the tight integration of all prior stages. This end-to-end capability is the hallmark of the most advanced, fully autonomous systems, such as The AI Scientist v1/v2 [Lu2024AIScientist, Yamada2025AIScientistV2], AI-Researcher [Tang2025AIResearcher], and DeepScientist [Weng2025DeepScientist].
@@ 2.2 Historical Evolution
  • Phase I: Foundational Modules (2022–2023). This initial phase established the foundational infrastructure for automating individual scientific workflow stages. Experimental preparation and execution received primary attention through benchmarks like DS-1000 [Lai2023DS1000], which formalized data science code generation tasks, and MLAgentBench [huang2023mlagentbench], which evaluated autonomous experimentation in machine learning contexts. Protocol design capabilities emerged through BioPlanner [BioPlanner2023], which translated natural language into executable biological protocols, establishing standards for biological experimental automation. Physical laboratory automation achieved a milestone through Coscientist [Boiko2023Coscientist], which integrated LLMs with robotic liquid handlers to demonstrate closed-loop chemical synthesis, validating the feasibility of AI-driven physical experimentation. Concurrently, early work in quantum computing experimentation [Cao2024QuantumAgentSDL] explored autonomous calibration and error correction, extending the automation paradigm to frontier physics domains. These systems collectively established modular capabilities across isolated workflow stages while revealing integration challenges.
@@
  • Phase II: Closed-Loop Integration (2024). The year 2024 witnessed a paradigm shift from isolated modules toward integrated workflows spanning multiple scientific stages. Literature synthesis capabilities matured through LitLLM [Agarwal2024LitLLM], which demonstrated citation-grounded knowledge extraction and narrative generation, establishing foundations for autonomous literature review. Hypothesis generation emerged as a distinct research thrust through SciAgents [Ghafarollahi2024SciAgents], which employed multi-agent collaboration for scientific ideation, and IdeaBench [guo2025ideabench], which formalized evaluation metrics for research question quality, novelty, and feasibility. The defining achievement of this phase was The AI Scientist v1 [Lu2024AIScientist], which unified literature review, idea generation, experimental execution, and manuscript writing into the first fully autonomous paper generation loop. This system demonstrated end-to-end research automation in machine learning domains, generating complete papers with minimal human intervention. Concurrently, specialized systems extended closed-loop reasoning to quantum computing [Cao2024QuantumAgentSDL] and materials science [Szymanski2023AutonomousLab], validating cross-domain generalization of integrated workflows.
@@
  • Phase III: Scalability, Impact, and Collaboration (2025–present). The contemporary frontier pursues three convergent objectives: scalability, scientific impact, and human-AI collaboration. Scalability and robustness advances through systems employing reinforcement learning and adaptive planning. HypER [vasu2025hyper] introduces provenance-aware hypothesis refinement through knowledge graph reasoning, while The AI Scientist v2 [Yamada2025AIScientistV2] implements agentic tree-search architectures enabling parallel exploration of research directions. DeepResearcher [Zheng2025DeepResearcher] extends scalability through reinforcement learning in web environments, training agents to navigate noisy, unstructured information. PiFlow [PiFlow2025] demonstrates principled scientific reasoning across domains through modular workflow orchestration. Scientific impact is targeted through systems designed for frontier discovery. ResearchBench [liu2025researchbench] establishes comprehensive evaluation protocols spanning ideation to publication, while DeepScientist [Weng2025DeepScientist] pursues goal-oriented, long-horizon research with explicit objectives to advance state-of-the-art. AI-Researcher [Tang2025AIResearcher] contributes Scientist-Bench for systematic assessment of autonomous research capabilities. Domain-specific advances continue through AutoLabs [Panapitiya2025AutoLabs] in chemistry, Curie [Curie2025] in ML experimentation, and SR-Scientist [SRScientist2025] in symbolic regression. Human-AI collaboration emerges as a critical thrust through frameworks like freephdlabor [Li2025FreephdLabor], which architectures research as interactive partnership between human strategists and AI executors, and AI co-scientist [gottweis2025towards], which formalizes collaborative research protocols.
@@ 3.1 Literature Review
    Stage 1: Corpus Acquisition and Layout-Aware Parsing. Ingests large-scale scientific repositories (arXiv, PubMed, Semantic Scholar) and transforms heterogeneous document formats into structured representations. Layout-aware parsing [lo2019s2orc] reconstructs logical document structure—sections, figures, tables, citations—normalizes metadata schemas, and builds citation graphs. This yields indexed corpora preserving textual content and structural semantics essential for downstream reasoning.
@@
    Stage 2: Retrieval and Re-Ranking. Optimizes information access through multi-stage ranking pipelines combining complementary strategies. Hybrid approaches leverage sparse lexical matching (BM25) and dense semantic encoders for conceptual similarity. Cross-encoder re-ranking refines candidates using contextualized scoring, while section-aware weighting prioritizes methodology sections. Systems like PaperQA [Lala2023PaperQA] integrate Retrieval-Augmented Generation, maintaining citation traceability.
@@
    Stage 3: Evidence Structuring and Representation Learning. Transforms unstructured textual evidence into machine-interpretable knowledge structures. Information extraction identifies entity mentions and relation tuples, knowledge graph induction maps findings to domain ontologies (UMLS, MeSH), and table synthesis [Deng2024T3] aggregates quantitative results. All transformations maintain explicit source provenance, enabling verification against original publications.
@@
    Stage 5: Grounded Narrative Synthesis and Novelty Analysis. Generates coherent research summaries while identifying unexplored directions. Retrieval-conditioned generation produces citation-grounded narratives anchored in source documents, contrastive novelty modeling identifies under-investigated areas, and iterative factuality verification [Agarwal2024LitLLM] detects inconsistencies. The output includes structured knowledge-gap maps highlighting promising opportunities.

### v5

@@ 1 Introduction
Artificial intelligence (AI) is undergoing a fundamental transformation in its role within scientific research. Historically deployed as a computational instrument for specific analytical tasks (such as pattern recognition, data mining, and predictive modeling [31, 3, 11]), AI is now emerging as an autonomous originator of scientific knowledge. This shift represents more than incremental technological progress; it embodies a qualitative leap from AI for Science, where AI augments human-defined workflows, to AI as Scientist, where AI systems independently conceptualize, execute, and communicate original research. The AI scientist paradigm pursues end-to-end autonomous discovery—executing the complete scientific workflow from hypothesis formulation through experimental execution to manuscript generation [16, 60, 66]. This transformation is enabled by three converging technological advances: (1) large language models with emergent reasoning capabilities [13]; (2) multi-agent architectures enabling complex task decomposition and collaboration [48, 14]; and (3) robotic experimentation platforms bridging digital reasoning with physical execution [24, 42]. Collectively, these advances mark a transition from automation—wherein AI supports predefined protocols—to autonomy—wherein AI independently designs, adapts, and validates investigative processes.
The intellectual lineage of autonomous scientific discovery extends to early symbolic AI systems [28, 48], which automated rediscovery of established physical laws through symbolic regression and rule-based reasoning. However, these pioneering efforts were constrained by brittle knowledge representations, limited generalization beyond carefully curated domains, and inability to interface with real-world experimentation. Contemporary AI Scientist systems transcend these limitations by integrating foundation models with closed-loop scientific reasoning (observe → hypothesize → experiment → analyze → publish), through agentic planning, principle-driven inference, and adaptive self-correction mechanisms. This integration enables systems to operate across heterogeneous knowledge modalities, navigate ambiguous problem spaces, and iteratively refine hypotheses based on empirical feedback. Representative implementations [36, 24, 44] demonstrate cross-domain capabilities spanning chemical synthesis [5], genomic network inference [1], and symbolic equation discovery [64], validating the generality of the autonomous research paradigm.
@@
Despite these advances, the field faces a critical challenge: conceptual fragmentation. The rapid proliferation of AI Scientist systems—evidenced by specialized frameworks achieving near-human performance [66, 55] and domain-specific benchmarks evaluating novelty, causality, and reproducibility [7, 33, 19]—has outpaced the development of unifying theoretical frameworks. Current research remains fragmented across three dimensions: (1) methodological silos, where systems excel at isolated stages (e.g., hypothesis generation [70, 59], literature synthesis [2], or experimental execution [24]) without addressing end-to-end integration; (2) domain-specific architectures, optimized for chemistry [5], biology [42], or physics with limited cross-domain generalization; and (3) evaluation heterogeneity, where benchmarks assess non-overlapping capabilities without establishing unified quality metrics. While existing surveys [16, 60, 37] provide valuable domain-specific or capability-focused reviews, no comprehensive synthesis systematically connects methodological stages, architectural patterns, and evaluation protocols across the full spectrum of AI Scientist research. This conceptual gap impedes cumulative progress, obscures transferable insights, and hinders the identification of fundamental research challenges versus domain-specific engineering issues.
This survey synthesizes research from 2022–2025 through a unified architectural framework (Figure 1). The horizontal axis defines six methodological stages: (1) Literature Review: knowledge structuring [2]; (2) Idea Generation: hypothesis formulation [59]; (3) Experimental Preparation: protocol design [26, 20]; (4) Experimental Execution: adaptive orchestration [42, 24, 43]; (5) Scientific Writing: multimodal composition [63]; and (6) Paper Generation: manuscript synthesis [55, 36]. The vertical axis stratifies four abstraction layers—Domains, Tasks, Architectures, Capabilities—forming a 4×6 categorization matrix. The timeline panel traces three evolutionary phases: Foundational Modules, Closed-Loop Integration, and Scalable Collaboration.
@@ 2 Background and Taxonomy
DS-1000 [26] 2023.04
Coscientist [5] 2023.06
BioPlanner [42] 2023.10
MLAgentBench [21] 2023.10
LitLLM [2] 2024.02
AI Scientist (v1) [36] 2024.08
SciAgents [14] 2024.09
IdeaBench [19] 2024.11
Quantum-Agent-SDL [6] 2024.12
HypER [59] 2025.01
AI Scientist (v2) [66] 2025.02
Curie [24] 2025.02
AI co-scientist [15] 2025.02
ResearchBench [33] 2025.03
DeepResearcher [68] 2025.04
AutoLabs [43] 2025.04
AI-Researcher [55] 2025.05
EXP-Bench [25] 2025.05
Agentic AutoSurvey [32] 2025.09
PiFlow [44] 2025.09
DeepScientist [61] 2025.09
SR-Scientist [64] 2025.10
Freephdlabor [29] 2025.10
@@ 2.1 Taxonomy of Research Categories
    Literature Review (Lit.). This foundational stage involves transforming unstructured scientific corpora into machine-interpretable knowledge. It encompasses techniques from large-scale information retrieval to the synthesis of research gaps. Systems like LitLLM [2] and HypER [59] focus on citation-grounded summarization and knowledge extraction, while advanced frameworks like SciAgents [14] and DeepResearcher [68] utilize web-scale interactions and graph reasoning to map existing knowledge and provide a robust foundation for downstream tasks.
@@
    Idea Generation (Idea). Building upon the structured knowledge from the prior stage, this phase automates hypothesis discovery and problem formulation. It leverages the creative and reasoning capabilities of LLMs to propose novel yet plausible research directions. This capability is explicitly evaluated by benchmarks like IdeaBench [19], and is a core component of both domain-specific systems like Coscientist [5] and advanced end-to-end frameworks like DeepScientist [61].
@@
    Experimental Preparation (Exp.). This crucial intermediate stage translates an abstract hypothesis into an executable plan. It includes tasks such as defining variables, selecting datasets, generating analysis code, and designing experimental protocols. This is a primary focus of data-science-oriented benchmarks like DS-1000 [26] and MLAgentBench [21], and is a key step in all integrated systems from The AI Scientist v1 [36] to freephdlabor [29].
@@
    Experimental Execution (Exec.). This stage involves the actual running of real or simulated experiments. It emphasizes the agent’s ability to interact with tools, control robotics, and adapt its plan based on real-time feedback. Milestones in this area include systems that orchestrate physical laboratory hardware, such as Coscientist [5] and Quantum-Agent-SDL [6]. Frameworks like Curie [24] and DeepResearcher [68] demonstrate this capability in simulated and real-world web environments, respectively.
@@
    Scientific Writing (Writ.). This stage focuses on the communication of scientific findings by transforming structured results into coherent, citation-grounded narratives. Capabilities range from section-aware summarization to data-to-text synthesis. This is a key feature in end-to-end systems like ResearchBench [33] and is central to human-in-the-loop frameworks like freephdlabor [29], where the AI drafts content for human review and refinement.
@@
    Paper Generation (Paper). Representing the culmination of the scientific workflow, this final stage synthesizes a full, publication-ready manuscript. This requires the tight integration of all prior stages. This end-to-end capability is the hallmark of the most advanced, fully autonomous systems, such as The AI Scientist v1/v2 [36, 66], AI-Researcher [55], and DeepScientist [61].
@@ 2.2 Historical Evolution
  • Phase I: Foundational Modules (2022–2023). This initial phase established the foundational infrastructure for automating individual scientific workflow stages. Experimental preparation and execution received primary attention through benchmarks like DS-1000 [26], which formalized data science code generation tasks, and MLAgentBench [21], which evaluated autonomous experimentation in machine learning contexts. Protocol design capabilities emerged through BioPlanner [42], which translated natural language into executable biological protocols, establishing standards for biological experimental automation. Physical laboratory automation achieved a milestone through Coscientist [5], which integrated LLMs with robotic liquid handlers to demonstrate closed-loop chemical synthesis, validating the feasibility of AI-driven physical experimentation. Concurrently, early work in quantum computing experimentation [6] explored autonomous calibration and error correction, extending the automation paradigm to frontier physics domains. These systems collectively established modular capabilities across isolated workflow stages while revealing integration challenges.
@@
  • Phase II: Closed-Loop Integration (2024). The year 2024 witnessed a paradigm shift from isolated modules toward integrated workflows spanning multiple scientific stages. Literature synthesis capabilities matured through LitLLM [2], which demonstrated citation-grounded knowledge extraction and narrative generation, establishing foundations for autonomous literature review. Hypothesis generation emerged as a distinct research thrust through SciAgents [14], which employed multi-agent collaboration for scientific ideation, and IdeaBench [19], which formalized evaluation metrics for research question quality, novelty, and feasibility. The defining achievement of this phase was The AI Scientist v1 [36], which unified literature review, idea generation, experimental execution, and manuscript writing into the first fully autonomous paper generation loop. This system demonstrated end-to-end research automation in machine learning domains, generating complete papers with minimal human intervention. Concurrently, specialized systems extended closed-loop reasoning to quantum computing [6] and materials science [54], validating cross-domain generalization of integrated workflows.
@@
  • Phase III: Scalability, Impact, and Collaboration (2025–present). The contemporary frontier pursues three convergent objectives: scalability, scientific impact, and human-AI collaboration. Scalability and robustness advances through systems employing reinforcement learning and adaptive planning. HypER [59] introduces provenance-aware hypothesis refinement through knowledge graph reasoning, while The AI Scientist v2 [66] implements agentic tree-search architectures enabling parallel exploration of research directions. DeepResearcher [68] extends scalability through reinforcement learning in web environments, training agents to navigate noisy, unstructured information. PiFlow [44] demonstrates principled scientific reasoning across domains through modular workflow orchestration. Scientific impact is targeted through systems designed for frontier discovery. ResearchBench [33] establishes comprehensive evaluation protocols spanning ideation to publication, while DeepScientist [61] pursues goal-oriented, long-horizon research with explicit objectives to advance state-of-the-art. AI-Researcher [55] contributes Scientist-Bench for systematic assessment of autonomous research capabilities. Domain-specific advances continue through AutoLabs [43] in chemistry, Curie [24] in ML experimentation, and SR-Scientist [64] in symbolic regression. Human-AI collaboration emerges as a critical thrust through frameworks like freephdlabor [29], which architectures research as interactive partnership between human strategists and AI executors, and AI co-scientist [15], which formalizes collaborative research protocols.
@@ 3.1 Literature Review
    Stage 1: Corpus Acquisition and Layout-Aware Parsing. Ingests large-scale scientific repositories (arXiv, PubMed, Semantic Scholar) and transforms heterogeneous document formats into structured representations. Layout-aware parsing [35] reconstructs logical document structure—sections, figures, tables, citations—normalizes metadata schemas, and builds citation graphs. This yields indexed corpora preserving textual content and structural semantics essential for downstream reasoning.
@@
    Stage 2: Retrieval and Re-Ranking. Optimizes information access through multi-stage ranking pipelines combining complementary strategies. Hybrid approaches leverage sparse lexical matching (BM25) and dense semantic encoders for conceptual similarity. Cross-encoder re-ranking refines candidates using contextualized scoring, while section-aware weighting prioritizes methodology sections. Systems like PaperQA [27] integrate Retrieval-Augmented Generation, maintaining citation traceability.
@@
    Stage 3: Evidence Structuring and Representation Learning. Transforms unstructured textual evidence into machine-interpretable knowledge structures. Information extraction identifies entity mentions and relation tuples, knowledge graph induction maps findings to domain ontologies (UMLS, MeSH), and table synthesis [10] aggregates quantitative results. All transformations maintain explicit source provenance, enabling verification against original publications.
@@
    Stage 5: Grounded Narrative Synthesis and Novelty Analysis. Generates coherent research summaries while identifying unexplored directions. Retrieval-conditioned generation produces citation-grounded narratives anchored in source documents, contrastive novelty modeling identifies under-investigated areas, and iterative factuality verification [2] detects inconsistencies. The output includes structured knowledge-gap maps highlighting promising opportunities.